# Interpreting CLIP with Hierarchical Sparse Autoencoders

### v1

@@ 4.4 Semantic Preservation Analysis
In Section 4.2, we only evaluated SAEs using $L_{0}$ for activation sparsity and EVR for reconstruction fidelity, however these metrics have limitations. $L_{0}$ only counts active neurons without assessing how well SAE representations align with original embeddings, and EVR focuses solely on distance reconstruction rather than semantic preservation. To address these limitations, we introduce additional metrics. Following (Yu et al., 2025), we adopt the CKNNA metric to assess how well SAE activations preserve the neighborhood structure of CLIP embeddings. We also evaluate semantic preservation through linear probing metrics; following (Gao et al., 2024; Lieberum et al., 2024), we use LP (KL) to measure prediction distribution alignment and LP (Acc) to compare classification accuracy. All metrics are defined in Section 4.1 and presented in Table 1. Our analysis reveals that while cosine similarity and FVU correlate well with linear probing metrics, the alignment metric demonstrates Matryoshka’s strength in preserving semantic structures.
@@ 4.8 Training Modality: Language and Vision
We evaluate how training modality affects MSAE performance by comparing models trained on text with the original image-trained models, validating both modality models across text and image domains in Table 2. While both variants perform best in their training domains, text-trained models achieve superior cross-modal performance, demonstrating stronger generalization capabilities. Moreover, text-trained models achieve higher sparsity on both modalities with no dead neurons, showing better utilization of learned features. These findings position text training as a preferred approach for multi-modal applications where balanced performance is desired. Future research could explore training SAEs on varying ratios of text and image data to optimize cross-modal performance or try to train crosscoders (Lindsey et al., 2024) on both modalities simultaneously.
We defer extended MSAE evaluations, also for ViT-B/16, to Appendix F.
@@ 5.1 Concept Naming
While self-supervised training of SAE enables learning up to $d$ monosemantic concepts, mapping these concepts to specific neurons remains non-trivial. Previous work used LLMs for identifying neuron-encoded concepts (Bills et al., 2023), but we adopt the more efficient method for CLIP-trained SAE proposed in (Rao et al., 2024), which leverages CLIP’s representation space. Our concept detection and validation methodology is detailed in Appendix A, with comprehensive results on valid concept counts across SAE models presented in Table 3. Figure 6 demonstrates the concept ’smile’ from MSAE RW (with 140 valid concepts detected) through its highest-activating text and image examples, confirming consistent concept presence across diverse inputs. Supplementary analysis of highly activated concept examples in Appendix G.1 showcase SAE’s ability to learn a wide range of concepts, from simple textures and colors to more complex ones like light (lights in darkness), countable concepts like trio (groups of three), and even nationality-related concepts like ireland or germany.
Figure 6: Visualization of the ‘smile’ concept through top-activating examples. Analysis of neuron no. 3552 in MSAE (RW) identified through automated interpretability as the concept smile with a similarity score of 0.59. Both text and image examples with the highest activation values strongly confirm the concept’s presence. We show additional such examples of valid concepts with their top-activating examples in Figure 19.
@@ 6 Conclusion
Limitations and future work. MSAE faces three limitations with clear paths for future improvement. The current implementation’s use of multiple decoder passes with different TopK activations introduces computational overhead, which could be addressed through optimized CUDA kernels enabling parallel processing of multiple granularities. While we demonstrated MSAE’s effectiveness using CLIP embeddings, it has great potential to explain hierarchical representations in other embedding spaces, such as SigLIP (Zhai et al., 2023) or modality-specific representations. Finally, since not all neurons correspond to simple concepts in our vocabulary, investigating complex semantic features through LLM-based interpretability methods could provide deeper insights into the learned hierarchical representations. MSAE has been proposed independently in concurrent work (Nabeshima, 2024; Bussmann et al., 2024a); however, we conduct comprehensive evaluations and focus on its concrete application to interpreting CLIP.
@@ Appendix B Implementation Details
We conducted experiments using CLIP ViT-L/14 (and ViT-B/16, reported later in the appendix) pre-trained on the CC3M dataset image training subset. Following (Bricken et al., 2023; Gao et al., 2024) and this blog [^1] https://transformer-circuits.pub/2024/april-update/index.html#training-saes, our SAE implementation uses unit-norm constraint on the decoder columns with untied encoder and decoder. We initialized $b_{\mathrm{pre}}$ as the geometric median of CLIP embeddings, $b_{\mathrm{enc}}$ with zeros, decoder with uniform Kaiming initialization (scaled $L_{2}$ norm to 0.1), and encoder as the decoder’s transpose. Gradient clipping was set to 1. For data preprocessing, we centralized embeddings per modality (Bhalla et al., 2024) and scaled by a constant to achieve $\mathbb{E}_{x\in\mathcal{X}}[\|x\|_{2}]=\sqrt{n}$. All models were trained for 30 epochs on a single NVIDIA A100 GPU with batch size 4096, except for the model with an expansion rate of 32, which was trained for 20 epochs. While MSAE and TopK showed dead neurons, we omitted revival strategies as only in TopK the number of dead neurons exceeded 1%.
@@ F.1 Sparsity–Fidelity Trade-off
Figure 11 presents an extended analysis of sparsity-fidelity trade-offs, including standard deviations and alternative reconstruction metric more tailored for the CLIP embeddings (cosine similarity). The results demonstrate MSAE’s superior stability across both modalities, particularly in text representations where only MSAE show stable and elevated results. Figure 12 strengthens our findings by showing MSAE superiority on different CLIP architecture (ViT-B/16).
Figure 11: Extended sparsity-fidelity trade-off analysis across modalities. Expanding on Figure 2, we compare ReLU SAE ($\lambda={0.001,0.003,0.03}$), TopK SAE ($k={32,64,128,256}$), and MSAE (UW, RW) using two reconstruction metrics: mean EVR fidelity (top) and mean cosine similarity (bottom). Results are shown for both image (left) and text (right) modalities, with standard deviation also reported for each metric, demonstrating MSAE’s consistent performance across modalities and metrics.
@@ F.5 Comprehensive Evaluation with CLIP ViT-L/14 and ViT-B/16 Architectures
We present an extensive quantitative comparison of SAE variants across CLIP ViT-L/14 and ViT-B/16 architectures. Our evaluation encompasses ReLU ($\lambda={0.03,0.003,0.001}$), TopK ($k={64,128,256}$), and MSAE (RW, UW) models, tested across three expansion rates (8, 16, 32) for both image and text modalities. For text modality and ViT-B/16 architecture, we omit LP (Acc) and LP (KL) metrics based on our findings in Section 4.4 that CS and FVU correlate strongly with linear probing metrics.
@@ F.5.1 Results for Image Modality
Table 6: CLIP ViT-L/14 SAE comparison at expansion rate 8. Extended evaluation from Table 1 with additional TopK ($k=128$) and ReLU ($\lambda=0.001$) variants on ImageNet-1k. Arrows indicate preferred metric direction, NDN values show training set dead neurons in parentheses.

### v2

@@ 4.4 Semantic Preservation Analysis
In Section 4.2, we only evaluated SAEs using $L_{0}$ for activation sparsity and EVR for reconstruction fidelity, however these metrics have limitations. $L_{0}$ only counts active neurons without assessing how well SAE representations align with original embeddings, and EVR focuses solely on distance reconstruction rather than semantic preservation. To address these limitations, we introduce additional metrics. Following (Yu et al., 2025), we adopt the CKNNA metric to assess how well SAE activations preserve the neighborhood structure of CLIP embeddings. We also evaluate semantic preservation through linear probing metrics (Gao et al., 2025; Lieberum et al., 2024), we use LP (KL) to measure prediction distribution alignment and LP (Acc) to compare classification accuracy. All metrics are defined in Section 4.1 and presented in Table 1. Our analysis reveals that while cosine similarity and FVU correlate well with linear probing metrics, the alignment metric demonstrates Matryoshka’s strength in preserving semantic structures.
@@ 4.8 Training Modality: Language and Vision
We evaluate how training modality affects MSAE performance by comparing models trained on text with the original image-trained models, validating both modality models across text and image domains in Table 2. While both variants perform best in their training domains, text-trained models achieve superior cross-modal performance, demonstrating stronger generalization capabilities. Moreover, text-trained models achieve higher sparsity on both modalities with no dead neurons, showing better utilization of learned features. These findings position text training as a preferred approach for multi-modal applications where balanced performance is desired. Future research could explore training SAEs on varying ratios of text and image data to optimize cross-modal performance or try to train crosscoders (Lindsey et al., 2024) on both modalities simultaneously. We defer extended MSAE evaluations, to Appendix F.
@@ 5.1 Concept Naming
While self-supervised training of SAE enables learning up to $d$ monosemantic concepts, mapping these concepts to specific neurons remains non-trivial. Previous work used LLMs for identifying neuron-encoded concepts (Bills et al., 2023), but we adopt the more efficient method for CLIP-trained SAE proposed in (Rao et al., 2024), which leverages CLIP’s representation space. Our concept detection and validation methodology is detailed in Appendix A, with comprehensive results on valid concept counts across SAE models presented in Table 3. Figure 8 demonstrates the highest-activating text and image examples for the best-matched feature concept ‘face’ across ReLU, TopK, and MSAE. The consistent concept presence across diverse inputs observed primarily in MSAE variants suggests that only Matryoshka-based methods were capable of learning this monosemantic feature. Supplementary analysis of highly activated concept examples in Appendix G.1 showcases SAE’s ability to learn a wide range of concepts, from simple textures and colors to more complex ones like light (lights in darkness), countable concepts like trio (groups of three), and even nationality-related concepts like ireland or germany.
@@ 5.3 Bias Validation on a Downstream Task
Figure 8: Comparison of top-activating examples for the ‘face’ concept across SAE methods. Through an automated interpretability process, we identified the best matching SAE neuron for the face concept, quantified by a similarity score and a validation status (Valid or Invalid). For this neuron, we then identified the highest activated examples in both text and image modalities across TopK, ReLU, and MSAE, each presented in variants optimized for either sparsity or reconstruction. Both text and image examples with the highest activation values strongly confirm. confirm the concept’s presence and demonstrate that only MSAE variants learned the concept of the face. We show additional examples of valid concepts with their top-activating examples in Figure 19.
@@ 6 Conclusion
Limitations and future work. MSAE faces three limitations with clear paths for future improvement. The current implementation’s use of multiple decoder passes with different TopK activations introduces computational overhead, which could be addressed through optimized CUDA kernels enabling parallel processing of multiple granularities. While we demonstrated MSAE’s effectiveness using CLIP embeddings, it has great potential to explain hierarchical representations in other embedding spaces, such as SigLIP (Zhai et al., 2023) or modality-specific representations. Finally, since not all neurons correspond to simple concepts in our vocabulary, investigating complex semantic features through LLM-based interpretability methods could provide deeper insights into the learned hierarchical representations. In concurrent independent work, Bussmann et al. (2025) propose an MSAE approach to interpreting language models.
@@
Acknowledgement
Work on this project was financially supported from the SONATA BIS grant 2019/34/E/ST6/00052 funded by Polish National Science Centre (NCN). We also thank the anonymous reviewers for their useful comments.
@@ Appendix B Implementation Details
We conducted experiments using CLIP ViT-L/14 (and ViT-B/16, reported later in the appendix) pre-trained on the CC3M dataset image training subset. Following (Bricken et al., 2023; Gao et al., 2025) and this blog (Conerly et al., 2024), our SAE implementation uses unit-norm constraint on the decoder columns with untied encoder and decoder. We initialized $b_{\mathrm{pre}}$ and $b_{\mathrm{enc}}$ with zeros, decoder with uniform Kaiming initialization (scaled $L_{2}$ norm to 0.1), and encoder as the decoder’s transpose. Gradient clipping was set to 1. For data preprocessing, we centralized embeddings per modality (Bhalla et al., 2024) and scaled by a constant to achieve $\mathbb{E}_{x\in\mathcal{X}}[\|x\|_{2}]=\sqrt{n}$. All models were trained for 30 epochs on a single NVIDIA A100 GPU with batch size 4096, except for the model with an expansion rate of 32, which was trained for 20 epochs. While MSAE and TopK showed dead neurons, we omitted revival strategies as only in TopK the number of dead neurons exceeded 1%.
@@ Appendix F Evaluating MSAE: Additional Results
Figure 11: Extended sparsity-fidelity trade-off analysis across modalities. Expanding on Figure 2, we compare ReLU SAE ($\lambda={0.03,0.01,0.003,0.001}$), TopK SAE ($k={32,64,128,256}$), BatchTopK SAE ($k={32,64,128,256}$), and MSAE (RW, UW) using two reconstruction metrics: mean EVR fidelity (top) and mean cosine similarity (bottom). Results are shown for both image (left) and text (right) modalities, with standard deviation also reported for each metric, demonstrating MSAE’s consistent performance across modalities and metrics.
@@ F.1 Sparsity–Fidelity Trade-off
Figure 11 presents an extended analysis of sparsity-fidelity trade-offs, including standard deviations and an alternative reconstruction metric tailored for CLIP embeddings (cosine similarity). The results demonstrate MSAE’s superior stability across both modalities, particularly in text representations where only MSAE models shows stable and elevated results. Furthermore, we observe that models learned with lower $k$ or higher sparsity regularization show high variance on the trained (image) modality, and this instability becomes even more pronounced on the text modality. Due to MSAE’s inherently low variance in the image modality, this similar low variance is preserved in the other modality, leading to consistently high and stable performance across both modalities. Figure 12 further strengthens our findings by showing MSAE’s superiority on a different CLIP architecture (ViT-B/16).
@@ F.5 Comprehensive Evaluation with CLIP ViT-L/14 and ViT-B/16 Architectures
We present an extensive quantitative comparison of SAE variants across CLIP ViT-L/14 and ViT-B/16 architectures. Our evaluation encompasses ReLU ($\lambda={0.03,0.01,0.003,0.001}$), TopK ($k={32,64,128,256,512}$), BatchTopK ($k={16,32,64,128,256}$), and MSAE (RW, UW) models, tested across three expansion rates (8, 16, 32) for both image and text modalities. For text modality and ViT-B/16 architecture, we omit LP (Acc) and LP (KL) metrics based on our findings in Section 4.4 that CS and FVU correlate strongly with linear probing metrics.
@@ F.5.1 Results for Image Modality
Table 6: CLIP ViT-L/14 SAE comparison at expansion rate 8. Extended evaluation from Table 1 with additional TopK, ReLU and BatchTopK variants on ImageNet-1k. Arrows indicate preferred metric direction, NDN values show training set dead neurons in parentheses.
@@
TopK ($k=512$) .$922_{\pm.015}.$$346_{\pm.442}.$$923_{\pm.058}.$ $56.1_{\pm 13.9}$$950_{\pm.218}.$$006_{\pm.003}.$$02.0$(1)010(1)0 ( 1 )
BatchTopK ($k=16$) .$698_{\pm.021}.$$371_{\pm.060}.$$798_{\pm.037}.$ $281_{\pm 326}$$836_{\pm.372}.$$698_{\pm.037}.$$02.0$(4278)042780(4278)0 ( 4278 )
BatchTopK ($k=32$) .$776_{\pm.019}.$$242_{\pm.034}.$$873_{\pm.020}.$ $113_{\pm 157}$$901_{\pm.299}.$$735_{\pm.004}.$$02.0$(3080)030800(3080)0 ( 3080 )
BatchTopK ($k=64$) .$877_{\pm.012}.$$162_{\pm.022}.$$917_{\pm.011}.$ $56.9_{\pm 85.8}$$931_{\pm.253}.$$769_{\pm.004}.$$02.0$(1477)014770(1477)0 ( 1477 )
BatchTopK ($k=128$) .$898_{\pm.010}.$$082_{\pm.012}.$$959_{\pm.006}.$ $23.3_{\pm 36.5}$$959_{\pm.197}.$$805_{\pm.004}.$$03.0$(539)05390(539)0 ( 539 )
BatchTopK ($k=256$) .$882_{\pm.005}.$$010_{\pm.005}.$$995_{\pm.002}.$ $2.42_{\pm 5.12}$$988_{\pm.108}.$$860_{\pm.003}.$002.002.002.002 $3(919)$
@@
TopK ($k=512$) .$964_{\pm.008}.$$336_{\pm.413}.$$926_{\pm.064}.$ $72.9_{\pm 17.8}$$944_{\pm.230}.$$007_{\pm.003}.$$02.0$(29)0290(29)0 ( 29 )
BatchTopK ($k=16$) .$669_{\pm.021}.$$404_{\pm.055}.$$786_{\pm.038}.$ $310_{\pm 353}$$829_{\pm.377}.$$705_{\pm.005}.$$01.0$(9859)098590(9859)0 ( 9859 )
BatchTopK ($k=32$) .$742_{\pm.021}.$$274_{\pm.031}.$$866_{\pm.020}.$ $122_{\pm 166}$$897_{\pm.304}.$$736_{\pm.005}.$$02.0$(8016)080160(8016)0 ( 8016 )
BatchTopK ($k=64$) .$880_{\pm.013}.$$167_{\pm.020}.$$916_{\pm.012}.$ $55.9_{\pm 85.0}$$932_{\pm.252}.$$759_{\pm.004}.$$02.0$(5113)051130(5113)0 ( 5113 )
BatchTopK ($k=128$) .$889_{\pm.011}.$$089_{\pm.012}.$$957_{\pm.006}.$ $24.6_{\pm 38.5}$$956_{\pm.204}.$$806_{\pm.003}.$002.002.002.002 $1(2967)$
BatchTopK ($k=256$) .$880_{\pm.009}.$$023_{\pm.006}.$$990_{\pm.003}.$ $4.12_{\pm 7.11}$$983_{\pm.131}.$$854_{\pm.003}.$002.002.002.002 $12(3558)$
@@
TopK ($k=512$) .$966_{\pm.004}.$$008_{\pm.005}.$$996_{\pm.002}.$ $1.30_{\pm 2.69}$$992_{\pm.092}.$$422_{\pm.018}.$002.002.002.002 $7822(22446)$
BatchTopK ($k=16$) .$638_{\pm.018}.$$462_{\pm.057}.$$775_{\pm.038}.$ $344.4_{\pm 377.9}$$825_{\pm.380}.$$735_{\pm.005}.$$01.0$(21631)0216310(21631)0 ( 21631 )
BatchTopK ($k=32$) .$756_{\pm.015}.$$329_{\pm.042}.$$855_{\pm.024}.$ $139.4_{\pm 186.8}$$894_{\pm.308}.$$747_{\pm.004}.$$01.0$(18965)0189650(18965)0 ( 18965 )
BatchTopK ($k=64$) .$880_{\pm.011}.$$190_{\pm.024}.$$911_{\pm.013}.$ $55.4_{\pm 84.7}$$934_{\pm.249}.$$764_{\pm.004}.$$02.0$(15035)0150350(15035)0 ( 15035 )
BatchTopK ($k=128$) .$869_{\pm.012}.$$129_{\pm.016}.$$947_{\pm.007}.$ $26.7_{\pm 42.2}$$956_{\pm.206}.$$794_{\pm.003}.$002.002.002.002 $1(11019)$
BatchTopK ($k=256$) .$837_{\pm.014}.$$069_{\pm.011}.$$979_{\pm.004}.$ $8.22_{\pm 13.4}$$976_{\pm.154}.$$851_{\pm.003}.$002.002.002.002 $12(11802)$
@@
TopK ($k=512$) .$882_{\pm.008}.$$058_{\pm.052}.$$972_{\pm.025}.$$005_{\pm.003}.$$03.0$(0)000(0)0 ( 0 )
BatchTopK ($k=16$) .$706_{\pm.023}.$$299_{\pm.050}.$$845_{\pm.030}.$$662_{\pm.005}.$$02.0$(2666)026660(2666)0 ( 2666 )
BatchTopK ($k=32$) .$810_{\pm.019}.$$196_{\pm.031}.$$900_{\pm.017}.$$695_{\pm.005}.$$03.0$(1763)017630(1763)0 ( 1763 )
BatchTopK ($k=64$) .$877_{\pm.013}.$$127_{\pm.020}.$$936_{\pm.010}.$$750_{\pm.004}.$$03.0$(830)08300(830)0 ( 830 )
BatchTopK ($k=128$) .$882_{\pm.010}.$$048_{\pm.008}.$$976_{\pm.004}.$$806_{\pm.003}.$$03.0$(387)03870(387)0 ( 387 )
BatchTopK ($k=256$) .$876_{\pm.003}.$$003_{\pm.001}.$$999_{\pm.001}.$$843_{\pm.003}.$003.003.003.003 $35(1766)$
@@
TopK ($k=512$) .$943_{\pm.015}.$$203_{\pm.312}.$$955_{\pm.047}.$$006_{\pm.003}.$003.003.003.003 1(1)111(1)1 ( 1 )
BatchTopK ($k=16$) .$658_{\pm.022}.$$349_{\pm.054}.$$837_{\pm.033}.$$670_{\pm.006}.$$02.0$(6478)064780(6478)0 ( 6478 )
BatchTopK ($k=32$) .$757_{\pm.022}.$$230_{\pm.035}.$$896_{\pm.019}.$$704_{\pm.005}.$$02.0$(5114)051140(5114)0 ( 5114 )
BatchTopK ($k=64$) .$853_{\pm.016}.$$135_{\pm.019}.$$938_{\pm.009}.$$741_{\pm.005}.$$03.0$(3145)031450(3145)0 ( 3145 )
BatchTopK ($k=128$) .$887_{\pm.010}.$$061_{\pm.011}.$$971_{\pm.005}.$$799_{\pm.004}.$$03.0$(1817)018170(1817)0 ( 1817 )
BatchTopK ($k=256$) .$876_{\pm.003}.$$003_{\pm.001}.$$999_{\pm.001}.$$843_{\pm.003}.$002.002.002.002 $40(4947)$
@@
TopK ($k=512$) .$967_{\pm.000}.$$001_{\pm.001}.$ $1.000_{\pm.001}$$016_{\pm.007}.$002.002.002.002 $5902(14864)$
BatchTopK ($k=16$) .$615_{\pm.018}.$$450_{\pm.060}.$$822_{\pm.037}.$$695_{\pm.005}.$$02.0$(14345)0143450(14345)0 ( 14345 )
BatchTopK ($k=32$) .$712_{\pm.021}.$$312_{\pm.043}.$$890_{\pm.019}.$$719_{\pm.005}.$$02.0$(12492)0124920(12492)0 ( 12492 )
BatchTopK ($k=64$) .$835_{\pm.017}.$$184_{\pm.028}.$$933_{\pm.010}.$$731_{\pm.005}.$002.002.002.002 $2(9645)$
BatchTopK ($k=128$) .$856_{\pm.013}.$$096_{\pm.015}.$$966_{\pm.005}.$$798_{\pm.003}.$$03.0$(6855)068550(6855)0 ( 6855 )
BatchTopK ($k=256$) .$844_{\pm.011}.$$040_{\pm.007}.$$991_{\pm.002}.$$827_{\pm.003}.$$02.0$(11638)0116380(11638)0 ( 11638 )
@@ F.5.2 Results for Text Modality
TopK ($k=512$) .$900_{\pm.026}.$$386_{\pm.376}.$$865_{\pm.101}.$$038_{\pm.006}.$$02.0$(1)010(1)0 ( 1 )
BatchTopK ($k=16$) .$585_{\pm.161}.$ $1.222_{\pm.929}$$416_{\pm.213}.$$031_{\pm.026}.$$02.0$(4278)042780(4278)0 ( 4278 )
BatchTopK ($k=32$) .$608_{\pm.221}.$$848_{\pm.277}.$$494_{\pm.232}.$$022_{\pm.012}.$$02.0$(3080)030800(3080)0 ( 3080 )
BatchTopK ($k=64$) .$712_{\pm.197}.$$662_{\pm.229}.$$581_{\pm.222}.$$019_{\pm.013}.$$02.0$(1477)014770(1477)0 ( 1477 )
BatchTopK ($k=128$) .$858_{\pm.038}.$$415_{\pm.180}.$$787_{\pm.102}.$$312_{\pm.009}.$$03.0$(539)05390(539)0 ( 539 )
BatchTopK ($k=256$) .$869_{\pm.026}.$$159_{\pm.175}.$$918_{\pm.108}.$$716_{\pm.004}.$002.002.002.002 $17(919)$
@@
TopK ($k=512$) .$950_{\pm.023}.$$435_{\pm.453}.$$855_{\pm.107}.$$073_{\pm.014}.$$02.0$(29)0290(29)0 ( 29 )
BatchTopK ($k=16$) .$600_{\pm.115}.$$917_{\pm.326}.$$450_{\pm.202}.$$032_{\pm.018}.$$01.0$(9859)098590(9859)0 ( 9859 )
BatchTopK ($k=32$) .$619_{\pm.181}.$$758_{\pm.209}.$$518_{\pm.232}.$$031_{\pm.021}.$$02.0$(8016)080160(8016)0 ( 8016 )
BatchTopK ($k=64$) .$706_{\pm.242}.$$687_{\pm.236}.$$559_{\pm.240}.$$029_{\pm.020}.$$02.0$(5113)051130(5113)0 ( 5113 )
BatchTopK ($k=128$) .$866_{\pm.034}.$$465_{\pm.230}.$$796_{\pm.081}.$$240_{\pm.016}.$002.002.002.002 $14(2967)$
BatchTopK ($k=256$) .$789_{\pm.189}.$$352_{\pm.371}.$$741_{\pm.334}.$$036_{\pm.004}.$002.002.002.002 $7(3558)$
@@
TopK ($k=512$) .$937_{\pm.062}.$$115_{\pm.131}.$$944_{\pm.069}.$$070_{\pm.034}.$002.002.002.002 $240(22446)$
BatchTopK ($k=16$) .$601_{\pm.079}.$$818_{\pm.209}.$$496_{\pm.191}.$$034_{\pm.024}.$$02.0$(21631)0216310(21631)0 ( 21631 )
BatchTopK ($k=32$) .$681_{\pm.143}.$$733_{\pm.235}.$$592_{\pm.190}.$$034_{\pm.020}.$$02.0$(21631)0216310(21631)0 ( 21631 )
BatchTopK ($k=64$) .$784_{\pm.178}.$$636_{\pm.225}.$$657_{\pm.181}.$$042_{\pm.026}.$$02.0$(18965)0189650(18965)0 ( 18965 )
BatchTopK ($k=128$) .$861_{\pm.016}.$$377_{\pm.159}.$$830_{\pm.055}.$$526_{\pm.012}.$002.002.002.002 $94(11019)$
BatchTopK ($k=256$) .$784_{\pm.139}.$$332_{\pm.333}.$$778_{\pm.288}.$$060_{\pm.007}.$$02.0$(11802)0118020(11802)0 ( 11802 )
@@
TopK ($k=512$) .$794_{\pm.174}.$$283_{\pm.282}.$$860_{\pm.154}.$$079_{\pm.021}.$$03.0$(0)000(0)0 ( 0 )
BatchTopK ($k=16$) .$602_{\pm.156}.$ $1.172_{\pm 1.110}$$461_{\pm.178}.$$018_{\pm.010}.$$02.0$(2666)026660(2666)0 ( 2666 )
BatchTopK ($k=32$) .$662_{\pm.213}.$$898_{\pm.510}.$$547_{\pm.176}.$$016_{\pm.008}.$$03.0$(1763)017630(1763)0 ( 1763 )
BatchTopK ($k=64$) .$716_{\pm.204}.$$656_{\pm.212}.$$637_{\pm.166}.$$029_{\pm.008}.$$03.0$(830)08300(830)0 ( 830 )
BatchTopK ($k=128$) .$715_{\pm.270}.$$654_{\pm.322}.$$662_{\pm.256}.$$033_{\pm.009}.$$03.0$(387)03870(387)0 ( 387 )
BatchTopK ($k=256$) .$774_{\pm.177}.$$317_{\pm.358}.$$776_{\pm.301}.$$103_{\pm.004}.$$03.0$(1766)017660(1766)0 ( 1766 )
@@
TopK ($k=512$) .$918_{\pm.051}.$$447_{\pm.804}.$$891_{\pm.089}.$$070_{\pm.009}.$$03.0$(1)010(1)0 ( 1 )
BatchTopK ($k=16$) .$592_{\pm.109}.$$940_{\pm.510}.$$480_{\pm.187}.$$018_{\pm.007}.$$02.0$(6478)064780(6478)0 ( 6478 )
BatchTopK ($k=32$) .$650_{\pm.174}.$$843_{\pm.428}.$$560_{\pm.182}.$$015_{\pm.007}.$$02.0$(5114)051140(5114)0 ( 5114 )
BatchTopK ($k=64$) .$734_{\pm.174}.$$649_{\pm.212}.$$654_{\pm.153}.$$022_{\pm.006}.$$03.0$(3145)031450(3145)0 ( 3145 )
BatchTopK ($k=128$) .$764_{\pm.218}.$$643_{\pm.310}.$$685_{\pm.216}.$$032_{\pm.004}.$$03.0$(1817)018170(1817)0 ( 1817 )
BatchTopK ($k=256$) .$799_{\pm.194}.$$337_{\pm.375}.$$755_{\pm.321}.$$061_{\pm.006}.$$02.0$(4947)049470(4947)0 ( 4947 )
@@
TopK ($k=512$) .$917_{\pm.109}.$$097_{\pm.172}.$$955_{\pm.087}.$$027_{\pm.016}.$002.002.002.002 $73(14864)$
BatchTopK ($k=16$) .$581_{\pm.066}.$$769_{\pm.191}.$$536_{\pm.157}.$$027_{\pm.012}.$$02.0$(14345)0143450(14345)0 ( 14345 )
BatchTopK ($k=32$) .$657_{\pm.113}.$$726_{\pm.284}.$$628_{\pm.136}.$$032_{\pm.020}.$$02.0$(12492)0124920(12492)0 ( 12492 )
BatchTopK ($k=64$) .$766_{\pm.126}.$$631_{\pm.227}.$$704_{\pm.113}.$$052_{\pm.024}.$$02.0$(9645)096450(9645)0 ( 9645 )
BatchTopK ($k=128$) .$783_{\pm.163}.$$628_{\pm.335}.$$727_{\pm.166}.$$046_{\pm.016}.$$03.0$(6855)068550(6855)0 ( 6855 )
BatchTopK ($k=256$) .$773_{\pm.172}.$$311_{\pm.339}.$$799_{\pm.270}.$$117_{\pm.006}.$$02.0$(11638)0116380(11638)0 ( 11638 )